OpenStudy (anonymous):

Use the five steps for problem solving to answer the following question. Please show all of your work. The area of the top of a triangular corner desk must be at least 7 sq. ft. If the base of the triangular desk top needs to be four feet, what should the height be? @Pre-Algebra

OpenStudy (anonymous):

Given : base = 4 feet area is 'at least' 7 sq feet To find : the height Formula: 1/2*b*h= area Solution: 1/2*4*h>= 7 2h>=7 h>=3.4

jhonyy9 (jhonyy9):

no the right value is h>= 7/2 so h>= 3,5